### v1

# ProtoSiTex: Learning Semi-Interpretable Prototypes for Multi-label Text Classification

Abstract.
The surge in user-generated reviews has amplified the need for interpretable models that can provide fine-grained insights. Existing prototype-based models offer intuitive explanations but typically operate at coarse granularity (sentence or document level) and fail to address the multi-label nature of real-world text classification. We propose ProtoSiTex, a semi-interpretable framework designed for fine-grained multi-label text classification. ProtoSiTex employs a dual-phase alternating training strategy: an unsupervised prototype discovery phase that learns semantically coherent and diverse prototypes, and a supervised classification phase that maps these prototypes to class labels. A hierarchical loss function enforces consistency across sub-sentence, sentence, and document levels, enhancing interpretability and alignment. Unlike prior approaches, ProtoSiTex captures overlapping and conflicting semantics using adaptive prototypes and multi-head attention. We also introduce a benchmark dataset of hotel reviews annotated at the sub-sentence level with multiple labels. Experiments on this dataset and two public benchmarks (binary and multi-class) show that ProtoSiTex achieves state-of-the-art performance while delivering faithful, human-aligned explanations, establishing it as a robust solution for semi-interpretable multi-label text classification.
Text Classification, Prototype Learning, Semi-interpretable Models
journalyear: 2025
1. Introduction
In recent years, online user reviews have become influential across domains like e-commerce, hospitality, and social media (Diaz et al., 2018). These reviews offer rich, unstructured insights into user preferences (Duan et al., 2008; Banerjee et al., 2017; Floyd et al., 2014). Extracting meaningful information from them is critical, and transformer-based language models (LMs) such as BERT (Devlin et al., 2019) have excelled in NLP tasks like sentiment analysis (Wankhade et al., 2022), question answering (Rogers et al., 2023), and text classification (Minaee et al., 2021). However, these models are typically opaque, limiting interpretability, an essential factor for trust, fairness, and compliance (Rudin, 2019; Castelvecchi, 2016). To address the limitations of black-box models, post-hoc explanation methods such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg et al., 2017) have been proposed. However, they often lack faithfulness and consistency, offering only approximate and sometimes inaccessible justifications (Crabbé et al., 2023; Alvarez-Melis et al., 2018), especially for non-technical users. This has led to growing interest in interpretable models that provide transparency by design (Molnar, 2020).
Prototype-Based Networks (PBNs) follow the paradigm of inherent interpretability, mimicking human reasoning via similarity to class-representative prototypes (Rosch, 1975). While effective in vision (Chen et al., 2019; Nauta et al., 2023) and recently adapted to NLP (Hong et al., 2023; Wen et al., 2025), most PBNs are limited to single-label, sentence- or document-level tasks. This limits their use in real-world, multi-label classification scenarios, e.g., “The view was breathtaking, but the staff was rude” expresses opposing sentiments, challenging coarse-grained prototype models (Sourati et al., 2024).
Efforts to improve interpretability in text classification often come at the expense of predictive strength. Transformer-based models capture rich semantic structures and typically outperform simpler methods in benchmarks (Li et al., 2022), yet their lack of transparency limits trust in sensitive applications. Prototype-driven approaches (Wei et al., 2024; Wen et al., 2025) provide clearer, case-based explanations that are easier for users to follow. However, they may sacrifice raw accuracy or struggle to scale to complex domains. This creates a persistent issue: optimizing solely for accuracy risks overlooking fairness and accountability, while prioritizing interpretability may reduce competitiveness in performance-driven evaluations. The issue becomes even more pronounced in multi-label classification (Pontiki et al., 2016; Schouten et al., 2015), where models must balance fine-grained predictive capability with explanations that remain faithful and comprehensible.
To address these limitations, we propose ProtoSiTex, a semi-interpretable prototype-based framework for fine-grained multi-label classification. ProtoSiTex effectively handles overlapping sentiments and multi-label scenarios via three key innovations: (i) adaptive prototype learning using multi-head attention for sub-sentence-level semantic alignment, (ii) a dual-phase alternating training strategy that combines unsupervised prototype discovery with supervised classification to improve generalization and reduce overfitting, and (iii) a hierarchical loss function that enforces consistency across sub-sentence, sentence, and document levels for structured and faithful predictions. We use the term semi-interpretable to denote models that provide faithful, prototype-based explanations for their predictions while still relying on latent neural mechanisms (e.g., attention, embeddings, hierarchical aggregation) that are not inherently transparent. Our key contributions are:
(i) A semi-interpretable classification framework using adaptive prototypes and multi-head attention: We introduce ProtoSiTex, a prototype-based model primarily for multi-label classification with interpretability. Unlike prior PBNs limited to sentence or document granularity, ProtoSiTex learns adaptive prototypes aligned with sub-sentence-level class patterns, enabling it to disentangle overlapping or conflicting sentiments. Leveraging multi-head attention, it captures diverse semantic interactions between text segments and prototypes, enhancing both alignment and interpretability.
(ii) Dual-phase alternating training strategy and hierarchical loss for robust learning and semantic consistency: To ensure generalizable and interpretable prototypes, we propose a dual-phase alternating learning strategy: (a) an unsupervised prototype discovery phase that promotes diversity and semantic coherence via clustering, and (b) a supervised classification phase that maps prototypes to multiple labels. This decoupling reduces overfitting by allowing periodic prototype refinement. We also introduce a hierarchical loss function to enforce consistency across sub-sentence, sentence, and document levels, enhancing structured reasoning, predictive accuracy, and explanation fidelity.
(iii) New dataset and extensive evaluation: We curate a hotel review dataset (HR) annotated at the sub-sentence-level with multi-label aspect tags, addressing the limitations of existing sentence or document-label benchmarks. To evaluate ProtoSiTex, we conducted extensive experiments on three datasets: (a) HR (multi-label classification), (b) IMDb (binary classification), and (c) TweetEval (multi-class classification). Across all datasets, ProtoSiTex consistently outperformed existing interpretable baselines and achieved performance on par with state-of-the-art black-box models, showing its effectiveness in terms of accuracy, robustness, and interoperability.
2. Related Work
Interpretable Text Classification: Deep models like CNNs, RNNs, and transformers have achieved strong performance across NLP tasks (Li et al., 2022), but their opaque nature hinders transparency. Post-hoc explanation methods, such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg et al., 2017), aim to interpret predictions by approximating local feature importance. However, these techniques operate independently of model training and often produce inconsistent or non-faithful rationales (Cesarini et al., 2024). This has spurred interest in interpretable models that learn to classify and explain simultaneously.
Prototype-based Networks: Prototype-Based Networks (PBNs) classify instances based on similarity to learned prototype vectors, offering human-aligned explanations via exemplar-based reasoning (Snell et al., 2017). Early works such as ProSeNet (Ming et al., 2019) and ClassVector (Yao et al., 2019a) introduced prototype learning for document-level, single-label tasks with limited interpretability. Subsequent methods like ProtoryNet (Hong et al., 2023) and ProtoLens (Wei et al., 2024) learned finer-grained prototypes at the sentence or subsentence level. GAProtoNet (Wen et al., 2025) integrated graph-attention with prototypes for relational reasoning. Yet, most PBNs focus on single-label classification, lacking the granularity and compositionality required for multi-label classification.
Multi-Label Text Classification: Multi-label text classification (MLTC) involves assigning multiple labels to a document, making it more challenging than single-label tasks due to inherent label dependencies and data imbalance (Liu et al., 2021). Early approaches relied primarily on problem transformation methods, such as Binary Relevance, Label power set, and Classifier Chains, as well as algorithm adaptation methods, including ML-SVMs and ML-kNN (Zhang et al., 2013). Models based on CNN and RNN improved text representation learning (Yang et al., 2016; Liu et al., 2017), while Graph Convolutional Networks were introduced to explicitly capture label correlations (Ma et al., 2021). More recently, Transformer-based architectures, including BERT (Devlin et al., 2019) have emerged as strong baselines. Despite these advances, existing methods either treat MLTC as a flat classification problem or rely on complex black-box architectures with limited interpretability. Furthermore, few datasets provide sub-sentence-level multi-label annotations, limiting empirical benchmarking.
Positioning: ProtoSiTex overcomes these limitations by unifying prototype-based interpretability with fine-grained multi-label classification. Unlike prior works, it learns adaptive prototypes aligned with sub-sentence semantics and leverages a dual-phase learning strategy, combining unsupervised prototype discovery and supervised hierarchical classification. This design enables structured predictions from sub-sentence to document level while providing reliable, human-understandable explanations.
3. Proposed Methodology
In this section, we present ProtoSiTex, our proposed method for semi-interpretable text classification.
Problem formulation: Given a document $\mathcal{D}$ composed of multiple sentences, our goal is to predict its associated class(es), accommodating both single-label and multi-label classification. ProtoSiTex performs document-level prediction while highlighting the specific content segments responsible for each label.
Preprocessing: In our framework, sub-sentence annotations, though optional, enhance fine-grained interpretability in multi-label tasks. Since manual labeling is costly, we use a few-shot LLM-based annotation engine (Dutta et al., 2025) to infer them from document-level labels automatically.
Each document $\mathcal{D}$ comprises $n$ sentences, represented as $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{n}\}$. Each sentence $\mathcal{S}_{i}$ is further segmented into $t_{i}$ sub-sentences, defined as $\mathcal{S}_{i}=\{s_{i1},s_{i2},\ldots,s_{it_{i}}\}$, where punctuation marks such as commas and semicolons are used as segmentation cues. The total number of sub-sentences in a document is denoted by $t=\sum_{i=1}^{n}t_{i}$. To assign fine-grained class labels, we apply the LLM-based engine to generate multi-label annotations at the sub-sentence level. These annotations are encoded in a binary label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$, where $l$ is the total number of classes. A matrix entry $c_{ij}=1$ indicates that the $i^{\text{th}}$ sub-sentence is associated with the $j^{\text{th}}$ label. For hierarchical supervision, we aggregate sub-sentence labels to form sentence-level labels, while document-level labels are directly provided. This bottom-up strategy ensures consistent multi-level annotation. To support structured supervision and hierarchical loss modeling, we construct two binary mapping matrices. The first, $\mathcal{M}^{1}\in\{0,1\}^{1\times n}$, captures the mapping from a document to its sentences. The second, $\mathcal{M}^{2}\in\{0,1\}^{n\times t}$, captures the mapping from each sentence to its associated sub-sentences. An entry $\mathcal{M}^{*}_{ij}=1$ indicates that sentence or sub-sentence $j$ belongs to sentence or document $i$, respectively.
Figure 1. Proposed solution architecture: ProtoSiTex
3.1. ProtoSiTex Architecture
The proposed framework, ProtoSiTex (Fig. 1), consists of three core stages: (i) encoding sub-sentences into high-dimensional semantic representations using a language model (LM)-based encoder, (ii) employing a dual-phase learning strategy that first discovers representative prototypes through unsupervised learning, then performs supervised classification guided by hierarchical supervision using document-level taxonomy, and (iii) performing inference through prototype-based reasoning applied to the learned embeddings. We detail each component below.
Sub-sentence Representation via LM-based Encoder: Effective text classification relies on context-aware representations. While traditional embeddings (e.g., Word2Vec, GloVe, FastText) are static, LM-based encoders offer dynamic, context-sensitive embeddings but lack interpretability. To balance expressiveness and transparency, we use an LM-based encoder to generate sub-sentence embeddings.
For each $s_{ij}\in\mathcal{S}_{i}$, the encoder outputs $e_{ij}\in\mathbb{R}^{d}$, producing a feature matrix $\mathcal{E}\in\mathbb{R}^{t\times d}$ over $t$ sub-sentences. This matrix feeds into the prototype discovery and classification modules, enabling both semantic precision and interpretability. We next describe the dual-phase learning.
Clustering-phase for Prototypes Learning: The first stage of our dual-phase learning aims to discover semantically meaningful and interpretable prototypes that facilitate multi-label classification. To achieve this, we introduce a learnable prototype layer coupled with a multi-head attention (MHA) mechanism, enabling diverse alignment between sub-sentence embeddings and prototypes.
We initialize $q$ prototypes $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{q}\};~\forall p_{i}\in\mathbb{R}^{d}$ using random initialization, which empirically outperforms K-Means, Xavier, and Kaiming. The value of $q$ is selected via grid search. Sub-sentence embeddings $\mathcal{E}\in\mathbb{R}^{t\times d}$ are projected into the prototype space using MHA, which learns flexible, context-aware alignments. For each attention head $i$, we compute query $\mathcal{Q}_{i}=\mathcal{P}\cdot\mathcal{W}^{{Q}_{i}}$, key $\mathcal{K}_{i}=\mathcal{E}\cdot\mathcal{W}^{{K}_{i}}$, and value $\mathcal{V}_{i}=\mathcal{E}\cdot\mathcal{W}^{{V}_{i}}$ matrices, where $\mathcal{W}^{Q_{i}},\mathcal{W}^{K_{i}},\mathcal{W}^{V_{i}}\in\mathbb{R}^{d\times d}$ are learnable parameters. Scaled dot-product attention computes the alignment:
(1) $\mathcal{H}_{i}=\sigma\left(\left(\mathcal{Q}_{i}\cdot\mathcal{K}_{i}^{\top}\right)/\sqrt{d}\right)\cdot\mathcal{V}_{i};\hskip 16.38895pt\mathcal{H}=\left(\mathcal{H}_{1}~\lVert~\mathcal{H}_{2}~\lVert...\lVert~\mathcal{H}_{h}\right)\cdot\mathcal{W}^{h}$
where, $\sigma$ denotes softmax function. The resulting matrix $\mathcal{H}\in\mathbb{R}^{q\times d}$ encodes prototype-aware representations of sub-sentences, which are used in the subsequent classification phase to associate prototypes with class labels.
Classification-phase for Learning Multi-granular Document Details: In the second phase, we train the classifier using prototype embeddings $\mathcal{P}$, sub-sentence embeddings $\mathcal{E}$, and prototype-aware representations $\mathcal{H}$. To enhance generalization, we apply multi-granular supervision without separate classifiers, enabling two-way explainability: from sub-sentences to prototypes and prototypes to labels. To map prototypes to class labels, we pass $\mathcal{H}$ through two dense layers with ReLU followed by a softmax function, producing prediction scores $\tilde{\mathcal{G}}_{1}\in\mathbb{R}^{q\times l}$:
(2) $\scriptsize\tilde{\mathcal{G}}_{1}={\sigma}\left(\mathcal{W}_{3}\cdot ReLU\left(\mathcal{W}_{2}\cdot ReLU\left(\mathcal{W}_{1}\cdot\mathcal{H}\right)\right)\right)$
Using $\tilde{\mathcal{G}}_{1}$, we derive sentence-level predictions $\tilde{\mathcal{G}}_{2}=\text{Clamp}\left(\mathcal{M}^{2}\cdot\mathcal{Y}^{\top}\cdot\tilde{\mathcal{G}}_{1}\right)$ $\in\mathbb{R}^{n\times l}$ by aggregating sub-sentence contributions. This is done via the binary prototype-to-subsentence assignment matrix $\mathcal{Y}$, computed by applying cosine similarity between prototypes and sub-sentences, and sentence-to-subsentence mapping $\mathcal{M}^{2}$. Finally, document-level predictions $\tilde{\mathcal{G}}_{3}=\text{Clamp}\left(\mathcal{M}^{1}\cdot\tilde{\mathcal{G}}_{2}\right)\in\mathbb{R}^{1\times l}$ are derived by aggregating sentence-level scores via document-sentence mapping $\mathcal{M}^{1}$. The $\text{Clamp}(\cdot)$ function restricts the elements of a matrix to the $[0,1]$ range by applying element-wise clipping, setting any negative value to 0 and any value above 1 to 1.
We generate dynamic ground-truths for training. Prototype-label associations $\mathcal{G}_{1}$ are computed by first measuring cosine similarity between prototypes and sub-sentence embeddings to form $\mathcal{X}=\mathcal{P}\cdot\mathcal{E}^{\top}/(\lVert\mathcal{P}\rVert\lVert\mathcal{E}^{\top}\rVert)$. This is binarized into $\mathcal{Y}\in\{0,1\}^{q\times t}$ as:
(3) $\scriptsize\mathcal{Y}_{ij}=\begin{cases}1,&\text{if }i=\arg\max_{k\in\{1~,\dots,~q\}}\mathcal{X}_{kj},\\ 0,&\text{otherwise}\end{cases}$
This binarized $\mathcal{Y}$ then multiplied with the sub-sentence label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$ to yield:
(4) $\scriptsize\mathcal{G}_{1}=\mathcal{Y}\cdot\mathcal{C};\qquad\mathcal{G}_{2}=\text{Clamp}(\mathcal{M}^{2}\cdot\mathcal{C})$
For sentence-level ground truth $\mathcal{G}_{2}$, we aggregate sub-sentence labels via $\mathcal{M}^{2}$. Document-level ground truth $\mathcal{G}_{3}\in\mathbb{R}^{1\times l}$ is directly taken from the dataset. It is worth noting that prototype and sentence-level labels are generated dynamically and depend on the current prototype configuration during the clustering phase.
3.2. Training Strategy of ProtoSiTex
We adopt an alternating dual-phase training strategy, where each epoch comprises $i_{1}$ clustering and $i_{2}$ classification iterations. Prototype parameters are updated in the clustering phase (with classifiers frozen), and vice versa in the classification phase. This decoupling stabilizes training, prevents early overfitting, and allows prototypes to evolve independently before being used for label prediction. We now describe the loss functions associated with each phase.
3.2.1. Loss functions for Clustering Phase
The clustering loss comprises two complementary components: proximity-based losses, which enforce alignment between sub-sentence embeddings and prototypes, and structure regularization losses, which ensure sparsity and diversity among the learned prototypes.
(i) Proximity-based Losses: These losses promote semantic alignment and coverage. The prototype matching loss ensures that each sub-sentence $\mathcal{E}_{i}$ is closely aligned to its nearest prototype $\mathcal{P}_{j^{*}_{i}}$, encouraging relevance. Simultaneously, the interpretability loss ensures each prototype remains active by encouraging proximity to at least one sub-sentence, thus achieving coverage and preventing prototype collapse. The joint formulation is given by:
(5) $\mathcal{L}_{prox}=-\frac{1}{t}\sum_{i=1}^{t}\left(\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j^{*}_{i}}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j^{*}_{i}}\rVert}\right)+\frac{1}{q}\sum_{j=1}^{q}\left(1-\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j^{*}_{i}}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j^{*}_{i}}\rVert}\right)$
were, $j^{*}_{i}$ is the index of the prototype closest to sub-sentence $\mathcal{E}_{i}$ based on cosine similarity.
(ii) Structure Regularization Losses: To enhance interpretability, we impose sparsity and diversity through regularization. The sparsity loss encourages selective activation, ensuring a limited set of sub-sentences influences each prototype. This reduces overlap and enhances clarity. The diversity loss, implemented via orthogonality regularization, pushes prototypes apart in embedding space, minimizing redundancy. The combined regularization is:
(6) $\mathcal{L}_{reg}=-\frac{1}{q}\sum_{j=1}^{q}\log\left(\sum_{i=1}^{t}\frac{\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{j}\right\rVert_{2})}{\sum_{k=1}^{q}\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{k}\right\rVert_{2})}+\epsilon\right)+\frac{1}{q(q-1)}\sum_{\begin{subarray}{c}i,j=1;\\ i\neq j\end{subarray}}^{q}\left(\frac{\mathcal{P}_{i}^{\top}\mathcal{P}_{j}}{\lVert\mathcal{P}_{i}\rVert\lVert\mathcal{P}_{j}\rVert}\right)^{2}$
The first term encourages peaked assignment distributions (sparsity), and the second term enforces pairwise dissimilarity (diversity) between prototypes.
To enforce semantic alignment with classes, we augment the clustering loss with the supervised classification loss $\mathcal{L}_{cs}$ (Eq. 8). The final clustering-phase loss ($\mathcal{L}_{cl}$) is:
(7) $\scriptsize\mathcal{L}_{cl}=\alpha_{1}~\mathcal{L}_{prox}+\alpha_{2}~\mathcal{L}_{reg}+\alpha_{3}~\mathcal{L}_{cs}$
Here, $\alpha_{1},\alpha_{2},\alpha_{3}$ are tunable weights summing to 1, controlling the influence of each component.
3.2.2. Loss functions for Classification Phase
The classification phase is designed to facilitate hierarchical supervision by capturing semantic patterns from fine-grained (sub-sentence) to coarse-grained (document) levels. To train this phase, we use three loss components: (i) a prototype-to-label interpretability loss that reinforces the semantic alignment between prototypes and classes, (ii) a sentence-wise supervision loss that captures mid-level contextual cues, and (iii) a document-level loss leveraging the provided annotations. Each loss component is computed using binary cross entropy (Paszke et al., 2019), and the combined loss is expressed as:
(8) $\scriptsize\mathcal{L}_{cs}=\lambda_{1}~\mathcal{L}_{1}(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1})+\lambda_{2}~\mathcal{L}_{2}(\mathcal{G}_{2},\tilde{\mathcal{G}}_{2})+\lambda_{3}~\mathcal{L}_{3}(\mathcal{G}_{3},\tilde{\mathcal{G}}_{3})$
Here, $\mathcal{G}_{i}$ and $\tilde{\mathcal{G}}_{i}$ (for $i=1,2,3$) represent ground-truth and predicted labels at prototype, sentence, and document levels. The weights $\lambda_{i}$ control each loss term, with $\sum_{i=1}^{3}\lambda_{i}=1$. Sub-sentence and prototype-level annotations are optional; when unavailable, $\lambda_{1}$ and $\lambda_{2}$ can be reduced or set to zero, enabling the model to rely solely on document-level labels and support flexible supervision.
This phase refines prototypes into interpretable classes, enhancing clustering and enabling generalizable, multi-granular predictions. Next, we describe the prototype-based inference process.
3.3. Prototype Projection and Label Classification
Each prototype $p_{i}\in\mathcal{P}$ is a high-dimensional vector residing in the same semantic space as sub-sentence embeddings. To ensure interpretability, we associate each prototype with its top-$K$ most similar sub-sentences from the training set, using cosine similarity. The top-$K$ nearest sub-sentences $\mathcal{T}_{K}(p_{i})$ are consolidated and passed to Gemini (Team et al., 2023) via an API to generate a concise summary reflecting their shared theme. This strategy ensures each prototype is grounded in semantically diverse and representative text snippets, offering robust explanations.
Given a test document $\mathcal{D}$, we segment it into sub-sentences $\{s_{ij}\}$, and encode them using the LM-based encoder to get $\mathcal{E}\in\mathbb{R}^{t\times d}$. These embeddings are projected to the prototype space via the trained multi-head attention module to produce prototype-aware representations $\mathcal{H}$, which are then passed through the proto-to-label classifier $\tilde{\mathcal{G}}_{1}$. Sentence- and document-level predictions ($\tilde{\mathcal{G}}_{2}$, $\tilde{\mathcal{G}}_{3}$) are obtained via hierarchical mappings using $\mathcal{M}^{2}$ and $\mathcal{M}^{1}$. The final class labels for the document are taken as $\tilde{\mathcal{G}}_{3}$. Simultaneously, we generate textual explanations by linking each sub-sentence to its closest prototype, and retrieving that prototype’s associated exemplars $\mathcal{T}_{K}(p_{i})$. This enables transparent reasoning behind predictions, making ProtoSiTex both accurate and interpretable.
ProtoSiTex is best described as semi-interpretable because it combines prototype-based explanations with deep neural components. On the one hand, its adaptive prototypes provide human-understandable evidence by aligning sub-sentences with semantically coherent exemplars, allowing users to see why particular text spans support specific class predictions. On the other hand, the model relies on latent mechanisms, multi-head attention, embedding transformations, and hierarchical aggregation across sub-sentence, sentence, and document levels, which are not directly transparent to human reasoning. As a result, the explanations are faithful and intuitive at the prototype level but do not offer full interpretability of the entire decision process. This balance between interpretability and accuracy justifies labeling the model as semi-interpretable rather than fully interpretable.
4. Experiments
All experiments were conducted on a Dell Precision 5860 Tower equipped with PyTorch 2.5.1, CUDA 12.2, an NVIDIA RTX A6000 GPU with 48 GB VRAM, and 256 GB RAM. We employed RoBERTa-Large (Liu et al., 2019) as the backbone for encoding subsentences in our framework. The HR dataset was split into training, validation, and test sets in a 7:2:17{:}2{:}1 ratio. For IMDb and TweetEVAL, we used the given training/ validation/ test sets. Each experiment was run five times, and the mean test performance is reported. Additional model configuration details are provided in the Appendix A.
4.1. Datasets
To evaluate the effectiveness of ProtoSiTex and compare it against state-of-the-art (SOTA) methods, we conducted comprehensive experiments on benchmark datasets IMDb (Maas et al., 2011), TweetEVAL (Barbieri et al., 2020), and our newly proposed Hotel Reviews (HR) dataset. These datasets encompass a diverse range of classification tasks, binary, multi-class, and multi-label, respectively.
(i) IMDb (Maas et al., 2011): Consisting of 50K movie reviews for binary sentiment classification (positive/ negative).
(ii) TweetEVAL (Barbieri et al., 2020): A multiclass dataset comprising 5052 tweets, with each tweet annotated with one of four emotion labels: anger, joy, optimism, or sadness.
(iii) Proposed Dataset: Hotel Reviews (HR): To support fine-grained, interpretable, multi-aspect classification in the hospitality domain, we introduce a novel dataset comprising hotel reviews, curated from tripadvisor.com. The reviews were collected via web scraping, ensuring a diverse coverage across locations, and customer sentiments. We identified $8$ key aspects based on extensive consultations with experts in travel management and linguistics (Fig. 2). These aspects were chosen to reflect both operational and experiential dimensions of hotel service quality. Each review may correspond to multiple aspect labels, reflecting the inherently multi-label nature of user opinions. To ensure high-quality annotations, we employed $10$ trained human annotators for each review. For every review, the annotators assigned one or more aspect labels as applicable, and highlighted the exact sub-sentence or phrase responsible for each aspect label, thereby providing ground-truth supervision for interpretability evaluation. Prior to annotation, all annotators underwent training. To ensure annotation consistency and reliability, we implemented a quality control mechanism using in-context learning via GPT-4o (Dutta et al., 2025). Annotator performance was iteratively assessed by comparing human-labeled segments, enabling correction and refinement through active feedback loops. The resulting dataset consists of 3002 reviews, annotated at the sub-sentence level with multi-label aspect tags and text spans responsible for each label, making it one of the first of its kind, publicly available resources to support interpretability-grounded multi-aspect classification at such fine granularity. Brief statistical details of HR are presented in Table 1.
Notably, sub-sentence level annotations are absent in the IMDb and TweetEVAL datasets.
Table 1. Statistics of proposed dataset: Hotel Reviews (HR)
Attributes Values No. of Hotels & Locations 275 & 70 No. of Reviews & Sentences 3002 & 26519 No. of Aspects 8 Sentence per review (Min & Max) 1 & 100 Sentence per review (Avg. & Std. Dev.) 8.83 & 7.47 Word per review (Avg.) 102.39
07507501,5001{,}5002,2502{,}2503,0003{,} [^0] Guest ExperienceLocation & SurroundingHousekeepingRoom & AmenitiesFoodConciergePersonalizationPricing & Value2,4172{,}4171,9551{,}9551,8501{,}8502,1872{,}1872,2852{,}2852,2552{,}255976976900900Number of Reviews
Figure 2. Distribution of eight aspects in HR dataset
4.2. Evaluation Metrics
We evaluate model performance using task-specific metrics. For binary and multi-class classification, we report Accuracy ($\mathcal{A}$) and F1-score ($\mathcal{F}$). Aspect-wise evaluation additionally includes Precision ($\mathcal{P}r$) and Recall ($\mathcal{R}$). For multi-label classification, we use macro-averaged F1-score ($\mathcal{F}_{m}$) and Balanced Accuracy ($\mathcal{BA}_{m}$). We also compute Hamming Loss ($\mathcal{HL}$) and ${\mathcal{HL}}^{c}$ to quantify label-wise prediction accuracy; ${\mathcal{HL}}^{c}=(1-\mathcal{HL})\times 100\%$.
Table 2. Comparison with baseline and SOTA approaches
[…]
(a) Refer to caption
(b) Refer to caption (c) Refer to caption
Figure 3. Qualitative analysis of ProtoSiTex on samples from (a) HR, (b) TweetEVAL (Barbieri et al., 2020), and (c) IMDb (Maas et al., 2011) datasets
4.3. Results & Comparative Study
To compare with ProtoSiTex, we chose 6 recent prototypical learning models: ProSeNet (Ming et al., 2019), ProtoTEX (Das et al., 2022), ProtoryNet (Hong et al., 2023), Proto-lm (Xie et al., 2023), GAProtoNet (Wen et al., 2025), and ProtoLens (Wei et al., 2025). In addition to this, we selected 9 black-box LM performances: ALBERT (Lan et al., 2020), DeBERTaV3 (He et al., 2023) BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2019), ELECTRA (Clark et al., 2020), XLNET (Yang et al., 2019), ModernBERT (Warner et al., 2025) RoBERTa and RoBERTa-Large (Liu et al., 2019) and 2 text classification methods: TextGCN (Yao et al., 2019b) and BertGCN (Lin et al., 2021), without incorporating interpretability.
In Table 2, we present the comparative study. This study reveals that existing prototype-based SOTA approaches often sacrifice performance for interpretability, leading to significantly lower results compared to baselines, particularly on challenging multi-label datasets such as HR. In contrast, ProtoSiTex successfully overcame this limitation by narrowing the performance gap with black-box LMs, while preserving inherent interpretability across varying dataset complexities. Specifically, ProtoSiTex achieved improvements of 0.04% and 2.85% in $\mathcal{A}$} on the IMDb and TweetEVAL datasets, respectively, and delivered a substantial 12.93% gain in $\mathcal{F}_{m}$} on HR, compared to the best prior prototype-based methods. These results across three diverse tasks highlight ProtoSiTex’s generalizability and applicability.
$\mathcal{F}$$\mathcal{BA}$${\mathcal{HL}}^{c}$75758080858590909595K-MeansKaiming NormXavier NormRandom Norm
$\mathcal{F}$$\mathcal{BA}$${\mathcal{HL}}^{c}$6060707080809090100100w/o ATProtoSiTex
$\mathcal{F}$$\mathcal{BA}$${\mathcal{HL}}^{c}$2020404060608080100100Euclidean Dist.Cosine Sim.ProtoSiTex
[…]
Figure 4. Performance analysis of ProtoSiTex on HR dataset: (a) with various prototype initializers; (b) without alternate training (w/o AT); (c) across different similarity measures; impact of the number of (d) prototypes, and (e) heads
[…]
Figure 5. Impact of hyperparameters $\alpha_{i}$ and $\lambda_{i}$
4.4. Qualitative Analysis of ProtoSiTex
To illustrate ProtoSiTex’s interpretability and prediction quality, we present qualitative examples from the HR, TweetEVAL, and IMDb datasets (Fig. 3). In each case, the model accurately predicts aspect or sentiment labels while grounding them in semantically coherent sub-sentence segments via learned prototypes. In the HR example, ProtoSiTex correctly identifies multi-label aspects (Room & Amenities, Food, Pricing & Value, etc.) from a single review, while providing text-grounded justifications. This highlights its ability to disentangle overlapping review aspects and attribute them to distinct prototype-based explanations. In the TweetEVAL sample, ProtoSiTex captures the emotion joy and connects it to meaningful prototypes reflecting positive sentiment, demonstrating its effectiveness in emotional reasoning with minimal text context. Similarly, in the IMDb example, ProtoSiTex predicts the positive label and supports it with a rich prototype cluster centered on artistic and narrative appreciation, confirming alignment with human sentiment. Overall, these examples demonstrate that ProtoSiTex offers not only accurate predictions but also transparent, faithful interpretations at the sub-sentence level, making it suitable for real-world applications requiring interpretability. More qualitative results are shown in Appendix B.
4.5. Impact of Model Components
Due to space constraints, we report the impact of the module only on the HR dataset, which is representative enough to illustrate the observed trends.
4.5.1. Impact of Different Prototype Initializers:
We evaluated four prototype initialization strategies: random, K-Means centroids, Kaiming Normal, and Xavier Normal. As shown in Fig. 4(a), random initialization consistently outperformed the others across $\mathcal{F}_{m}$, $\mathcal{BA}_{m}$, and ${\mathcal{HL}}^{c}$. This performance gain is attributed to the increased diversity introduced early in the training, which helps avoid premature convergence and overfitting.
4.5.2. Impact of Alternate Training:
ProtoSiTex adopts a dual-phase training strategy, alternating between unsupervised prototype learning and supervised classification. To evaluate its efficacy, we compared model variants trained with and without alternation. As shown in Fig. 4(b), alternating training significantly improved performance by enabling the model to discover semantically meaningful prototype clusters and align them with downstream classification objectives. In contrast, without alternation, the model struggled to learn task-specific prototypes, lowering performance.
4.5.3. Impact of Similarity Function:
We evaluated three similarity metrics: MHA, cosine similarity, and Euclidean distance, for associating subsentence embeddings with prototypes. As presented in Fig. 4(c), MHA of ProtoSiTex achieved superior $\mathcal{F}_{m}$, $\mathcal{BA}_{m}$, and ${\mathcal{HL}}^{c}$ scores by capturing richer embedding relationships, leading to more discriminative and interpretable prototype assignments.
4.6. Impact of Hyperparameters
ProtoSiTex includes several key hyperparameters: the number of prototypes ($q$), number of heads ($h$), loss weights for clustering ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$), and classification ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$). To systematically investigate the influence of these hyperparameters on model performance, we conducted a series of experiments on the HR dataset. As shown in Fig. 4(d), increasing $q$ initially boosts performance by capturing diverse semantic concepts, but excessive values lead to redundancy and over-smoothing. A similar trend is observed for $h$ in Fig. 4(e). Furthermore, Fig. 5 shows that assigning $\alpha_{1}=0.01$, $\alpha_{2}=0.98$, $\alpha_{3}=0.01$, $\lambda_{1}=0.6$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.2$ yields optimal results. This multi-level supervision prevented the model from overfitting to a single level of granularity, supporting both local and global interpretability and accuracy.
[…]
Figure 6. Aspect-wise performance of ProtoSiTex on HR
(a) Refer to caption
(b) Refer to caption (c) Refer to caption
Figure 7. Misprediction examples of ProtoSiTex on (a) HR, (b) TweetEVAL (Barbieri et al., 2020), and (c) IMDb (Maas et al., 2011) datasets
4.7. Aspect-wise Performance Analysis:
Fig. 6 presents a detailed aspect-wise evaluation of ProtoSiTex on the HR dataset, offering insights into the model’s performance across individual aspect categories. Notably, the aspects Guest Experience, Food, and Concierge not only appear the most frequently in HR but also exhibit strong predictive performance. In contrast, Pricing & Value and Personalization are the least represented. Despite the challenges posed by class imbalance, ProtoSiTex shows a strong ability to identify the Pricing & Value aspect, achieving $\mathcal{F}$ and $\mathcal{A}$ scores comparable to those of the more frequent categories. However, Personalization remains the most challenging aspect, particularly in terms of $\mathcal{R}$, as the model often fails to capture all relevant instances. This difficulty is likely due to both its low frequency in the dataset and its inherently subjective nature. Unlike more explicit aspects such as Food or Location & Surrounding, Personalization is often expressed through subtle, indirect, and context-dependent language, making it more difficult to detect.
4.8. Misprediction Analysis
To better understand the behavior and limitations of ProtoSiTex, we conducted a detailed analysis of mispredictions. This analysis highlights recurring patterns that lead to incorrect predictions. The most notable cases of mispredictions are discussed below:
(i) Boundary Ambiguity in Prototype: In multi-label classification, ProtoSiTex occasionally exhibits boundary ambiguity among prototypes, where semantically related class representations capture overlapping evidence. This ambiguity leads to confusion in distinguishing co-occurring categories, resulting in the assignment of only a single dominant label instead of multiple relevant ones. For example, in Fig. 7(a), the review “Spacious room with balcony with amazing view of the valley and the fall”. The actual labels are Room & Amenities and Location & Surrounding. However, ProtoSiTex predicts only Room & Amenities. Here, Room & Amenities dominates and captures shared contextual characteristics such as “spacious room”, “balcony”, overshadowing the equally valid cues of “view of the valley” that should activate Location & Surrounding.
(ii) Lexical cue dominance: Sometimes the mispredictions in ProtoSiTex were caused by lexical bias, the tendency to rely excessively on high-intensity words when making predictions. Words such as “angry”, “worst”, or “best” often dominate the model’s decision process, even when their contextual meaning is softened/inverted by surrounding text. For instance, in Fig. 7(b) the review “i love ur angry comments”, the actual emotion is affectionate teasing or joy, yet the ProtoSiTex incorrectly predicts anger due to the presence of the word “angry”.
(iii) Absence of Fine-Grained Annotation: This observation highlights an important takeaway: even when only coarse document-level supervision is available, prototype-based explanations remain highly useful for error analysis and model trust. By showing how individual sub-sentences align with prototypes tied to different target classes, the model offers interpretable evidence that pinpoints why a misclassification occurred, rather than treating it as a black-box error. From a technical perspective, the misclassification arises due to aggregation bias, while sub-sentence–prototype mappings are accurate, the hierarchical aggregation process (from prototype-level activations to the final document-level decision) can overweight strong local cues and misjudge the overall label. In practice, this means that even when the document-level prediction is wrong, the prototype-level reasoning still faithfully explains the decision path, helping to uncover issues such as prototype imbalance, lack of discourse-level cues (e.g., concessions or sarcasm), or dataset limitations tied to coarse supervision. Thus, beyond accuracy metrics, our approach demonstrates that interpretable prototype assignments provide actionable insights for debugging and model refinement. As illustrated in Fig. 7(c), the identified prototypes, explanations, and associated class mappings are correct, even though the final document-level prediction was misclassified.
Conclusion
We proposed ProtoSiTex, a semi-interpretable framework for fine-grained, multi-label text classification. By combining adaptive prototype learning, dual-phase alternating training, and hierarchical supervision, ProtoSiTex delivers accurate and transparent predictions across multiple levels of granularity. To support evaluation, we introduced the Hotel Reviews (HR) dataset with sub-sentence-level multi-aspect annotations. Extensive experiments on HR, along with the publicly available IMDb and TweetEVAL datasets, show that ProtoSiTex matches or exceeds baseline and state-of-the-art models, while providing faithful, human-aligned interpretations. Future directions include extending ProtoSiTex to cross-domain and multilingual settings, incorporating multimodal cues, and integrating large language models for interactive, user-centric explanations.
Appendix A Hyperparameters Configuration
Table 3. Brief details of hyperparameters
Attribute Value
Optimizer AdamW
Learning rate (Clustering-phase) $2\times 10^{-5}$
Learning rate (Classification-phase) $2\times 10^{-5}$
Batch size 64
LM encoder (RoBERTa-Large): Feature dimension 1024
No. of dense layer 2
Dimension of dense layer 256 & 64
Clustering loss parameters ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ ) [0-1]
Classification loss parameters ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$) [0-1]
No. of head [1-32]
No. of prototype [4-64]
Table 3 presents the hyperparameters engaged in ProtoSiTex. To obtain high-dimensional, contextually rich text embeddings, we incorporate a language model, called RoBERTa (Liu et al., 2019). Specifically, it is a robustly optimized BERT variant, which we employed in its large-weight version pretrained on large English-language internet corpora using the masked language modeling objective. RoBERTa-large remains tunable within our framework. We maintain a consistent dropout rate of 0.1 across all trainable weights. To avoid overfitting, early stopping with a patience of three epochs based on validation performance is applied. Additionally, we set a maximum training limit of 100 epochs. Each epoch includes 10 iterations of the clustering phase and 100 iterations of the classification phase, ensuring a balanced optimization of both phases.
Appendix B Qualitative Analysis
While we presented a qualitative analysis in the main paper, this section provides a more detailed examination, offering finer-grained insights into the interpretability and effectiveness of ProtoSiTex. Tables 4 show textual representation of learned prototypes along with their corresponding classes, derived after training on all datasets (Hotel Reviews (HR), IMDb (Maas et al., 2011), and TweetEVAL (Barbieri et al., 2020)). Our analysis reveals that the learned prototypes are well-distributed across each dataset and span all labels. Additionally, we show qualitative results for samples from HR, IMDb, and TweetEVAL in Fig.s 8.
B.1. Analysis on Hotel Reviews (HR)
Fig. 8(a) demonstrates the interpretability of ProtoSiTex on multi-aspect text. Here, the hotel review referencing honeymoon context directly supports the prototype of “Fostering Memorable Celebrations and Immersive Activities”, while expressions such as “hotel has the best view to offer” closely match the prototype “Location: Scenic and Strategically Advantageous Settings”. ProtoSiTex also provides rich prototype justification for additional aspects and successfully predicts the review into Concierge, Personalization, Food, Location & Surrounding, Guest Experience aspects.
B.2. Analysis on IMDb
In Fig. 8(b), ProtoSiTex classifies the input text as positive, recognizing its emphasis on how this film version outshines all previous adaptations. The model captures the sentiment of novelty and strength in storytelling, with a strong alignment to relevant prototypes, demonstrating both the interpretability and effectiveness of ProtoSiTex. Similarly, ProtoSiTex predicts negative for the sample presented in Fig. 8(c).
B.3. Analysis on TweetEval
Fig. 8(d-g) illustrate the two-way interpretability of ProtoSiTex on TweetEVAL. Specifically, Fig. 8(d), the input text reflects deep emotional pain rooted in past trauma, conveying feelings of depression and grief. This aligns closely with the prototype capturing themes of “Existential dread and loss”, showcasing ProtoSiTex’s ability to contextualize emotional content through semantically coherent prototypes.
(a)Refer to caption
(b) Refer to caption (c) Refer to caption
(d) Refer to caption (e) Refer to caption
(f) Refer to caption (g) Refer to caption
Figure 8. ProtoSiTex qualitative analysis on samples from (a) HR, (b)-(c) IMDb (Maas et al., 2011), and (d)-(g) TweetEVAL (Barbieri et al., 2020) datasets
Table 4. Textual representative of learned prototypes on All Datasets
Sl. Prototype Texts Labels HR Dataset (Ours) 1 Arrival logistics, personalization, and service recovery Concierge, Personalization, Guest Experience 2 Personalization and navigating external factors Personalization, Guest Experience 3 Inconsistent functionality and safety concerns Room & Amenities, Guest Experience 4 Complimentary room upgrades and enhanced stay benefits Personalization, Room & Amenities 5 Exceptional guest experience through proactive service and unexpected delights Guest Experience, Concierge 6 Room quality and prime location versus maintenance and staff issues Room & Amenities, Location & Surrounding, Concierge 7 Housekeeping: excellence in service marred by pest issues Housekeeping 8 Room features, accessibility, and guest privacy concerns. Room & Amenities, Guest Experience 9 Varied impressions of hotel property and service Guest Experience, Room & Amenities 10 State of facilities and guest offerings Room & Amenities 11 Assessing cost-effectiveness and perceived worth Pricing & Value 12 Well-maintained facilities and friendly service. Room & Amenities, Concierge 13 Exceptional staff and room quality Room & Amenities, Housekeeping 14 Location: proximity, environmental noise, and scenic aspects Location & Surrounding 15 Navigating operational shortcomings and external challenges Guest Experience, Concierge 16 Cleanliness and quality: rooms, amenities, and breakfast Housekeeping, Room & Amenities, Food 17 Inconsistent guest experience and room and amenities quality Room & Amenities, Guest Experience 18 Varied staff performance and property conditions with notable location and dining Guest Experience, Room & Amenities, Food, Location & Surrounding 19 Variable quality across guest experience and property aspects Guest Experience, Room & Amenities 20 Balancing location advantages with pricing concerns Location & Surrounding, Pricing & Value 21 Overpriced with inadequate service and amenities Pricing & Value, Room & Amenities 22 Recurring room and amenity malfunctions and service deficiencies Room & Amenities, Housekeeping 23 Financial dealings and perceived value Pricing & Value 24 Inconsistent value for money with unexpected additional costs Pricing & Value 25 Spacious and family-friendly Room & Amenities 26 Quality and utility of facilities and in-room provisions Room & Amenities 27 Varied room types, upgrades, and amenity performance Room & Amenities 28 Diverse guest insights across hotel operations, location ambiance, and service delivery Guest Experience, Location & Surrounding, Concierge 29 Location: scenic and strategically advantageous settings Location & Surrounding 30 Contrasting guest experience and perceived value Guest Experience, Pricing & Value 31 Service and billing management Concierge, Pricing & Value 32 Elevated stays through thoughtful amenities and diverse culinary offerings Room & Amenities, Food, Guest Experience 33 Discrepant billing, payment obstacles, and perceived poor value Pricing & Value 34 Outstanding hospitality, culinary delights, and comfort. Food, Concierge 35 Varying quality in food, facilities, and service Food 36 Focus on renovation and comfort upgrades Room & Amenities 37 Handling of booking, billing, and diverse guest requests Concierge 38 Property condition, aesthetic appeal, and service quality Room & Amenities 39 Prominent on-site dining and accommodation features Food, Room & Amenities 40 Diverse dining experiences and efficient guest services Food , Concierge 41 Fostering memorable celebrations and immersive activities Guest Experience 42 Mixed perceptions of room condition, amenities, and service quality relative to cost Room & Amenities, Guest Experience, Pricing & Value 43 Location: a decisive factor for guests Location & Surrounding 44 Hotel’s location and on-site environment: accessibility, amenities, and challenges Location & Surrounding, Room & Amenities 45 Food & beverage: value, variety, and pricing transparency Food, Pricing & Value 46 Staff engagement and support Concierge 47 Personalized attention and special arrangements Personalization, Concierge 48 Modern amenities and varied room conditions with notable maintenance and service lapses Room & Amenities, Housekeeping, Guest Experience IMDB Dataset (Maas et al., 2011) 1 Unveiling the depth of human emotion through diverse narratives Positive 2 Discerning appraisals of cinematic,theatrical performances, adaptations Positive 3 Celebrating unconventional and engrossing storytelling Positive 4 Potent narratives of human struggle and unflinching reality Positive 5 Diverse perspectives and defended convictions Positive 6 Cinema of intense debate and baffling choices Positive 7 Memorable, albeit flawed, cinematic experiences Negative 8 Films with commendable elements, yet falling short of their potential Negative TweetEVAL Dataset (Barbieri et al., 2020) 1 Irritation and frustration Anger 2 Anger and disgust Anger 3 Irritation and disrespect triggered outbursts Anger 4 Humorous and offended outbursts Anger 5 Contentment and reflection Joy 6 Cuddling and positivity promote well-being Joy 7 Gratitude and joyful appreciation Joy 8 Regret, resilience, and hope amidst adversity Optimism 9 Minor discomforts and empathetic support Optimism 10 Gloomy discomfort Sadness 11 Disappointment and discouragement Sadness 12 Existential dread and loss Sadness