### v2

# ProtoSiTex: Learning Semi-Interpretable Prototypes for Multi-label Text Classification

Abstract
The surge in user-generated reviews has amplified the need for interpretable models that can provide fine-grained insights. Existing prototype-based models offer intuitive explanations but typically operate at coarse granularity (sentence or document level) and fail to address the multi-label nature of real-world text classification. We propose ProtoSiTex, a semi-interpretable framework designed for fine-grained multi-label text classification. ProtoSiTex employs a dual-phase alternate training strategy: an unsupervised prototype discovery phase that learns semantically coherent and diverse prototypes, and a supervised classification phase that maps these prototypes to class labels. A hierarchical loss function enforces consistency across subsentence, sentence, and document levels, enhancing interpretability and alignment. Unlike prior approaches, ProtoSiTex captures overlapping and conflicting semantics using adaptive prototypes and multi-head attention. We also introduce a benchmark dataset of hotel reviews annotated at the subsentence level with multiple labels. Experiments on this dataset and two public benchmarks (binary and multi-class) show that ProtoSiTex achieves state-of-the-art performance while delivering faithful, human-aligned explanations, establishing it as a robust solution for semi-interpretable multi-label text classification.
I Introduction
In recent years, online user reviews have become influential across domains like e-commerce, hospitality, and social media [review_helpfulness]. These reviews offer rich, unstructured insights into user preferences [online_review, online_review_1, online_review_2]. Extracting meaningful information from reviews is critical, and transformer-based language models (LMs) [devlin-etal-2019-bert] have excelled in NLP tasks like sentiment analysis [sentiment_survey], question answering [QA_survey], and text classification [TC_survey]. However, these models are typically opaque, limiting interpretability, an essential factor for trust, fairness, and compliance [interpretability_1, interpretability_2]. To address the limitations of black-box models, post-hoc explanation methods such as LIME [LIME] and SHAP [SHAP] have been proposed. However, they often lack faithfulness and consistency, offering only approximate and sometimes inaccessible justifications [interpretability_3, interpretability_4], especially for non-technical users. This has led to growing interest in interpretable models that provide transparency by design [interpretability_5].
Prototype-Based Networks (PBNs) follow the paradigm of inherent interpretability, mimicking human reasoning via similarity to class-representative prototypes [cognitive]. While effective in vision [proto_img1, proto_img2] and recently adapted to NLP [10.5555/3648699.3648963-ProtoryNet, wen-etal-2025-gaprotonet], most PBNs are limited to single-label, sentence- or document-level tasks. This limits their use in real-world, multi-label classification scenarios, e.g., “The view was breathtaking, but the staff was rude” expresses opposing sentiments, challenging coarse-grained prototype models [sourati-etal-2024-robust].
Efforts to improve interpretability in text classification often come at the expense of predictive strength. Transformer-based models capture rich semantic structures and typically outperform simpler methods in benchmarks [10.1145/3495162-text-class-survey], yet their lack of transparency limits trust in sensitive applications. Prototype-driven approaches [protolens, wen-etal-2025-gaprotonet] provide clearer, case-based explanations that are easier for users to follow. However, they may sacrifice raw accuracy or struggle to scale to complex domains. This creates a persistent issue: optimizing solely for accuracy risks overlooking fairness and accountability, while prioritizing interpretability may reduce competitiveness in performance-driven evaluations. The issue becomes even more pronounced in multi-label classification [pontiki2016semeval, schouten2015survey], where models must balance fine-grained predictive capability with explanations that remain faithful and comprehensible.
To address these limitations, we propose ProtoSiTex, a semi-interpretable prototype-based framework for fine-grained multi-label classification. ProtoSiTex effectively handles overlapping sentiments and multi-label scenarios via three key innovations: (a) adaptive prototype learning using multi-head attention for subsentence-level semantic alignment, (b) a dual-phase alternate training strategy that combines unsupervised prototype discovery with supervised classification to improve generalization and reduce overfitting, and (c) a hierarchical loss function that enforces consistency across subsentence, sentence, and document levels for structured and faithful predictions. We use the term semi-interpretable to denote models that provide faithful, prototype-based explanations for their predictions while still relying on latent neural mechanisms (e.g., attention, embeddings, hierarchical aggregation) that are not inherently transparent. Our key contributions are:
(i) A semi-interpretable classification framework using adaptive prototypes and multi-head attention: We introduce ProtoSiTex, a prototype-based model primarily for multi-label classification with interpretability. Unlike prior PBNs limited to sentence or document granularity, ProtoSiTex learns adaptive prototypes aligned with subsentence-level class patterns, enabling it to disentangle overlapping or conflicting sentiments. Leveraging multi-head attention, it captures diverse semantic interactions between text segments and prototypes, enhancing both alignment and interpretability.
(ii) Dual-phase alternate training strategy and hierarchical loss for robust learning and semantic consistency: To ensure generalizable and interpretable prototypes, we propose a dual-phase alternate learning strategy: (a) an unsupervised prototype discovery phase to promote diversity and semantic coherence via clustering, and (b) a supervised classification phase that maps prototypes to multiple labels. This decoupling reduces overfitting by allowing periodic prototype refinement. We also introduce a hierarchical loss function to enforce consistency across subsentence, sentence, and document levels, enhancing structured reasoning, predictive accuracy, and explanation fidelity.
(iii) New dataset and extensive evaluation: We curate a hotel reviews dataset (HR) annotated at the subsentence-level with multi-label aspect tags, addressing the limitations of existing sentence or document-label benchmarks. To evaluate ProtoSiTex, we conducted extensive experiments on three datasets: (a) HR (multi-label classification), (b) IMDb (binary classification), and (c) TweetEval (multi-class classification). Across all datasets, ProtoSiTex consistently outperformed existing interpretable baselines and achieved performance on par with state-of-the-art black-box models, showing its effectiveness in terms of accuracy, robustness, and interoperability.
The rest of the paper is organized as follows: Section II briefly reviews the literature, Section III presents our proposed approach, Section IV details the experiments, and Section V concludes with directions for future work.
II Related Work
This section briefly surveys related work and positions our study within the literature.
Interpretable Text Classification: Deep models like CNNs, RNNs, and transformers have achieved strong performance across NLP tasks [10.1145/3495162-text-class-survey], but their opaque nature hinders transparency. Post-hoc explanation methods, such as LIME [LIME] and SHAP [SHAP], aim to interpret predictions by approximating local feature importance. However, these techniques operate independently of model training and often produce inconsistent or non-faithful rationales [cesarini2024explainable-posthoc-limitation]. This has spurred interest in interpretable models that learn to classify and explain simultaneously.
Prototype-Based Networks (PBNs): PBNs classify instances based on similarity to learned prototype vectors, offering human-aligned explanations via exemplar-based reasoning [snell2017prototypical_CV]. Early works such as ProSeNet [10.1145/3292500.3330908-ProsNet] and ClassVector [10.1145/3318299.3318307-ClassVector] introduced prototype learning for document-level, single-label tasks with limited interpretability. Subsequent methods like ProtoryNet [10.5555/3648699.3648963-ProtoryNet] and ProtoLens [protolens] learned finer-grained prototypes at the sentence or subsentence level. GAProtoNet [wen-etal-2025-gaprotonet] integrated graph-attention with prototypes for relational reasoning. Yet, most PBNs focus on single-label classification, lacking the granularity and compositionality required for multi-label classification.
Multi-Label Text Classification (MLTC): MLTC involves assigning multiple labels to a document, making it more challenging than single-label tasks due to inherent label dependencies and data imbalance [multi_label_tpami]. Early approaches relied primarily on problem transformation methods, such as binary relevance, label power set, and classifier chains, as well as algorithm adaptation methods, including multi-label SVMs and KNNs [multi_label_methods]. Models based on CNNs and RNNs improved text representation learning [multi_label_RNN, multi_label_cnn], while GCNs were introduced to explicitly capture label correlations [multi_label_gcn]. More recently, transformer-based architectures, including BERT [devlin-etal-2019-bert] have emerged as strong baselines. Despite these advances, existing methods either treat MLTC as a flat classification problem or rely on complex black-box architectures with limited interpretability. Furthermore, there is hardly any dataset that provides subsentence-level multi-label annotations, limiting empirical benchmarking.
Positioning: ProtoSiTex overcomes these limitations by unifying prototype-based interpretability with fine-grained multi-label classification. Unlike prior works, it learns adaptive prototypes aligned with subsentence semantics and leverages a dual-phase learning strategy, combining unsupervised prototype discovery and supervised hierarchical classification. This design enables structured predictions from subsentence to document level while providing reliable, human-understandable explanations.
III Proposed Methodology
In this section, we present ProtoSiTex, our proposed method for semi-interpretable text classification.
Figure 1: Proposed solution architecture: ProtoSiTex
III-A Problem Formulation
Given a document $\mathcal{D}$ composed of multiple sentences, our goal is to predict its associated class(es), accommodating both single-label and multi-label classification. ProtoSiTex performs document-level prediction while highlighting the specific content segments responsible for each label.
Preprocessing: In our framework, subsentence annotations, though optional, enhance fine-grained interpretability in multi-label tasks. Since manual labeling is costly, we use a few-shot LLM-based annotation engine [ARTICLE_AAAI_2025] to infer them from document-level labels automatically.
Each document $\mathcal{D}$ comprises $n$ sentences, represented as $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{n}\}$. Each sentence $\mathcal{S}_{i}$ is further segmented into $t_{i}$ subsentences, defined as $\mathcal{S}_{i}=\{s_{i1},s_{i2},\ldots,s_{it_{i}}\}$, where punctuation marks such as commas and semicolons are used as segmentation cues. The total number of subsentences in a document is denoted by $t=\sum_{i=1}^{n}t_{i}$. To assign fine-grained class labels, we apply the LLM-based engine to generate multi-label annotations at the subsentence level. These annotations are encoded in a binary label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$, where $l$ is the total number of classes. A matrix entry $c_{ij}=1$ indicates that the $i^{\text{th}}$ subsentence is associated with the $j^{\text{th}}$ label. For hierarchical supervision, we aggregate subsentence labels to form sentence-level labels, while document-level labels are directly provided. This bottom-up strategy ensures consistent multi-level annotation. To support structured supervision and hierarchical loss modeling, we construct two binary mapping matrices. The first, $\mathcal{M}^{1}\in\{0,1\}^{1\times n}$, captures the mapping from a document to its sentences. The second, $\mathcal{M}^{2}\in\{0,1\}^{n\times t}$, captures the mapping from each sentence to its associated subsentences. An entry $\mathcal{M}^{2}_{ij}=1$ indicates that sentence or subsentence $j$ belongs to sentence or document $i$, respectively.
III-B ProtoSiTex Architecture
The proposed framework, ProtoSiTex (Fig. 1), consists of three core stages: (i) encoding subsentences into high-dimensional semantic representations using a language model (LM)-based encoder, (ii) employing a dual-phase learning strategy that first discovers representative prototypes through unsupervised learning, then performs supervised classification guided by hierarchical supervision using document-level taxonomy, and (iii) performing inference through prototype-based reasoning applied to the learned embeddings. We detail each component below.
III-B1 Subsentence Representation via LM-based Encoder
Effective text classification relies on context-aware representations. While traditional embeddings (e.g., Word2Vec, GloVe, FastText) are static, LM-based encoders offer dynamic, context-sensitive embeddings but lack interpretability. To balance expressiveness and transparency, we use an LM-based encoder to generate subsentence embeddings.
For each $s_{ij}\in\mathcal{S}_{i}$, the encoder outputs $e_{ij}\in\mathbb{R}^{d}$, producing a feature matrix $\mathcal{E}\in\mathbb{R}^{t\times d}$ over $t$ subsentences. This matrix feeds into the prototype discovery and classification modules, enabling both semantic precision and interpretability. We next describe the dual-phase learning.
III-B2 Dual Phase Learning
The clustering and classification phases aim to learn prototypes and capture multi-granular document representations, respectively, as elaborated below.
Clustering-phase for Prototypes Learning
The first stage of our dual-phase learning aims to discover semantically meaningful and interpretable prototypes that facilitate multi-label classification. To achieve this, we introduce a learnable prototype layer coupled with a multi-head attention (MHA) mechanism, enabling diverse alignment between subsentence embeddings and prototypes.
We initialize $q$ prototypes $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{q}\};\penalty 10000\ \forall p_{i}\in\mathbb{R}^{d}$ using random initialization, which empirically outperforms K-means, Xavier, and Kaiming. The value of $q$ is selected via grid search.
Subsentence embeddings $\mathcal{E}\in\mathbb{R}^{t\times d}$ are projected into the prototype space using MHA, which learns flexible, context-aware alignments. For each attention head $i$, we compute query $\mathcal{Q}_{i}=\mathcal{P}{\cdot}\mathcal{W}^{{Q}_{i}}$, key $\mathcal{K}_{i}=\mathcal{E}{\cdot}\mathcal{W}^{{K}_{i}}$, and value $\mathcal{V}_{i}=\mathcal{E}{\cdot}\mathcal{W}^{{V}_{i}}$ matrices, where $\mathcal{W}^{Q_{i}},\mathcal{W}^{K_{i}},\mathcal{W}^{V_{i}}\in\mathbb{R}^{d\times d}$ are learnable parameters. Scaled dot-product attention computes the alignment:
$\mathcal{H}_{i}=\sigma\left(\left(\mathcal{Q}_{i}{\cdot}\mathcal{K}_{i}^{\top}\right)/\sqrt{d}\right){\cdot}\mathcal{V}_{i};\hskip 16.38895pt\mathcal{H}=\left(\mathcal{H}_{1}\penalty 10000\ \lVert\penalty 10000\ \mathcal{H}_{2}\penalty 10000\ \lVert...\lVert\penalty 10000\ \mathcal{H}_{h}\right){\cdot}\mathcal{W}^{h}$ (1)
where, $\sigma$ denotes softmax function. The resulting matrix $\mathcal{H}\in\mathbb{R}^{q\times d}$ encodes prototype-aware representations of subsentences, which are used in the subsequent classification phase to associate prototypes with class labels.
Classification-phase for Learning Multi-granular Document Details
In the second phase, we train the classifier using prototype embeddings $\mathcal{P}$, subsentence embeddings $\mathcal{E}$, and prototype-aware representations $\mathcal{H}$. To enhance generalization, we apply multi-granular supervision without separate classifiers, enabling two-way explainability: from subsentences to prototypes and prototypes to class-labels. To map prototypes to labels, we pass $\mathcal{H}$ through two dense layers with ReLU followed by a softmax function, producing prediction scores $\tilde{\mathcal{G}}_{1}\in\mathbb{R}^{q\times l}$:
$\footnotesize\tilde{\mathcal{G}}_{1}={\sigma}\left(\mathcal{W}_{3}{\cdot}ReLU\left(\mathcal{W}_{2}{\cdot}ReLU\left(\mathcal{W}_{1}{\cdot}\mathcal{H}\right)\right)\right)$ (2)
Using $\tilde{\mathcal{G}}_{1}$, we derive sentence-level predictions $\tilde{\mathcal{G}}_{2}=f_{clamp}\left(\mathcal{M}^{2}{\cdot}\mathcal{Y}^{\top}{\cdot}\tilde{\mathcal{G}}_{1}\right)$ $\in\mathbb{R}^{n\times l}$ by aggregating subsentence contributions. This is done via the binary prototype-to-subsentence assignment matrix $\mathcal{Y}$, computed by applying cosine similarity between prototypes and subsentences, and sentence-to-subsentence mapping $\mathcal{M}^{2}$.
Finally, document-level predictions $\tilde{\mathcal{G}}_{3}=f_{clamp}\left(\mathcal{M}^{1}{\cdot}\tilde{\mathcal{G}}_{2}\right)\in\mathbb{R}^{1\times l}$ are derived by aggregating sentence-level scores via document-sentence mapping $\mathcal{M}^{1}$.
The $f_{clamp}(\cdot)$ function restricts the elements of a matrix to the $[0,1]$ range by applying element-wise clipping, setting any negative value to 0 and any value above 1 to 1.
We generate dynamic ground-truths for training. Prototype-label associations $\mathcal{G}_{1}$ are computed by first measuring cosine similarity between prototypes and subsentence embeddings to form $\mathcal{X}=\mathcal{P}{\cdot}\mathcal{E}^{\top}/(\lVert\mathcal{P}\rVert\lVert\mathcal{E}^{\top}\rVert)\in\mathbb{R}^{q\times t}$ that is binarized into $\mathcal{Y}$ as:
$\footnotesize\mathcal{Y}_{ij}=\begin{cases}1,&\text{if }i=\arg\max_{k\in\{1\penalty 10000\ ,\dots,\penalty 10000\ q\}}\mathcal{X}_{kj},\\ 0,&\text{otherwise}\end{cases}$ (3)
The binarized $\mathcal{Y}\in\{0,1\}^{q\times t}$ is then multiplied with the subsentence label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$ to yield:
$\footnotesize\mathcal{G}_{1}=\mathcal{Y}\cdot\mathcal{C};\qquad\mathcal{G}_{2}=f_{clamp}(\mathcal{M}^{2}\cdot\mathcal{C})$ (4)
For sentence-level ground truth $\mathcal{G}_{2}$, we aggregate subsentence labels via $\mathcal{M}^{2}$. Document-level ground truth $\mathcal{G}_{3}\in\mathbb{R}^{1\times l}$ is directly taken from the dataset. It is worth noting that prototype and sentence-level labels are generated dynamically and depend on the current prototype configuration during the clustering phase.
III-B3 Prototype Projection and Label Classification
Each prototype ${\mathcal{P}}_{i}\in\mathcal{P}$ is a high-dimensional vector residing in the same semantic space as subsentence embeddings. To ensure interpretability, we associate each prototype with its top-$K$ most similar subsentences from the training set, using cosine similarity. The top-$K$ nearest subsentences $\mathcal{T}_{K}({\mathcal{P}}_{i})$ are consolidated and passed to Gemini Pro 2.5 [gemini2023] to generate a concise summary reflecting their shared theme. This strategy ensures each prototype is grounded in semantically diverse and representative text snippets, offering robust explanations.
Given a test document $\mathcal{D}$, we segment it into subsentences, and encode them using the LM-based encoder to get $\mathcal{E}\in\mathbb{R}^{t\times d}$. These embeddings are projected to the prototype space via the trained multi-head attention module to produce prototype-aware representations $\mathcal{H}$, which are then passed through the proto-to-label classifier $\tilde{\mathcal{G}}_{1}$. Sentence- and document-level predictions ($\tilde{\mathcal{G}}_{2}$, $\tilde{\mathcal{G}}_{3}$) are obtained via hierarchical mappings using $\mathcal{M}^{2}$ and $\mathcal{M}^{1}$. The final predicted class labels for the document are taken as $\tilde{\mathcal{G}}_{3}$. Simultaneously, we generate textual explanations by linking each subsentence to its closest prototype, and retrieving that prototype’s associated exemplars $\mathcal{T}_{K}({\mathcal{P}}_{i})$. This enables transparent reasoning behind predictions, making ProtoSiTex both accurate and interpretable.
ProtoSiTex is best described as semi-interpretable because it combines prototype-based explanations with deep neural components. On the one hand, its adaptive prototypes provide human-understandable evidence by aligning subsentences with semantically coherent exemplars, allowing users to see why particular text spans support specific class predictions. On the other hand, the model relies on latent mechanisms, multi-head attention, embedding transformations, and hierarchical aggregation across subsentence, sentence, and document levels, which are not directly transparent to human reasoning. As a result, the explanations are faithful and intuitive at the prototype level but do not offer full interpretability of the entire decision process. This balance between interpretability and accuracy justifies labeling the model as semi-interpretable rather than fully interpretable.
III-C Training Strategy of ProtoSiTex
We adopt an alternate dual-phase training strategy, where each epoch comprises $i_{1}$ clustering and $i_{2}$ classification iterations. Prototype parameters are updated in the clustering phase (with classifiers frozen), and vice versa in the classification phase. This decoupling stabilizes training, prevents early overfitting, and allows prototypes to evolve independently before being used for label prediction. We now describe the loss functions associated with each phase.
III-C1 Loss Functions for Classification Phase
The classification phase is designed to facilitate hierarchical supervision by capturing semantic patterns from fine-grained (subsentence) to coarse-grained (document) levels. To train this phase, we use three loss components: (i) a prototype-to-label interpretability loss that reinforces the semantic alignment between prototypes and classes, (ii) a sentence-wise supervision loss that captures mid-level contextual cues, and (iii) a document-level loss leveraging the provided annotations. Each loss component is computed using binary cross entropy [PyTorch], and the combined loss is expressed as:
$\footnotesize\mathcal{L}_{cs}=\lambda_{1}\penalty 10000\ \mathcal{L}_{1}(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1})+\lambda_{2}\penalty 10000\ \mathcal{L}_{2}(\mathcal{G}_{2},\tilde{\mathcal{G}}_{2})+\lambda_{3}\penalty 10000\ \mathcal{L}_{3}(\mathcal{G}_{3},\tilde{\mathcal{G}}_{3})$ (5)
where, $\mathcal{G}_{i}$ and $\tilde{\mathcal{G}}_{i}$ (for $i=1,2,3$) represent ground-truth and predicted labels at prototype, sentence, and document levels. The weights $\lambda_{i}$ control each loss term, with $\sum_{i=1}^{3}\lambda_{i}=1$. Subsentence and prototype-level annotations are optional; when unavailable, $\lambda_{1}$ and $\lambda_{2}$ can be reduced or set to zero, enabling the model to rely solely on document-level labels and support flexible supervision.
III-C2 Loss Functions for Clustering Phase
The clustering loss comprises two complementary components: proximity-based losses that enforce alignment between subsentence embeddings and prototypes, and structure regularization losses that ensure sparsity and diversity among the learned prototypes.
Proximity-based Losses
These losses promote semantic alignment and coverage. The prototype matching loss ensures that each subsentence $\mathcal{E}_{i}$ is closely aligned to its nearest prototype $\mathcal{P}_{j^{*}_{i}}$, encouraging relevance. Simultaneously, the interpretability loss ensures each prototype remains active by encouraging proximity to at least one subsentence, thus achieving coverage and preventing prototype collapse. The joint formulation is given by:
$\mathcal{L}_{prox}=-\frac{1}{b}\sum_{i=1}^{b}\left(\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j^{*}_{i}}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j^{*}_{i}}\rVert}\right)+\frac{1}{q}\sum_{j=1}^{q}\left(1-\max_{i=1}^{b}\left(\frac{\mathcal{E}_{i}^{\top}\mathcal{P}_{j}}{\lVert\mathcal{E}_{i}\rVert\lVert\mathcal{P}_{j}\rVert}\right)\right)$ (6)
where $j^{*}_{i}$ is the index of the prototype closest to subsentence $\mathcal{E}_{i}$ based on cosine similarity, and $b$ denotes the total number of subsentences across all reviews in a mini-batch.
Structure Regularization Losses
To enhance interpretability, we impose sparsity and diversity through regularization. The sparsity loss encourages selective activation, ensuring a limited set of subsentences influences each prototype. This reduces overlap and enhances clarity. The diversity loss, implemented via orthogonality regularization, pushes prototypes apart in embedding space, minimizing redundancy. The combined regularization is:
$\mathcal{L}_{reg}=-\frac{1}{q}\sum_{j=1}^{q}\log\left(\sum_{i=1}^{b}\frac{\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{j}\right\rVert_{2})}{\sum_{k=1}^{q}\exp(-\left\lVert\mathcal{E}_{i}-\mathcal{P}_{k}\right\rVert_{2})}+\epsilon\right)$ (7)
$+\frac{1}{q(q-1)}\sum_{\begin{subarray}{c}i,j=1;\\ i\neq j\end{subarray}}^{q}\left(\frac{\mathcal{P}_{i}^{\top}\mathcal{P}_{j}}{\lVert\mathcal{P}_{i}\rVert\lVert\mathcal{P}_{j}\rVert}\right)^{2}$
where, the first term encourages peaked assignment distributions (sparsity), and the second term enforces pairwise dissimilarity (diversity) between prototypes.
To enforce semantic alignment with classes, we augment the clustering loss with the supervised classification loss $\mathcal{L}_{cs}$. The final clustering-phase loss ($\mathcal{L}_{cr}$) is:
$\footnotesize\mathcal{L}_{cr}=\alpha_{1}\penalty 10000\ \mathcal{L}_{prox}+\alpha_{2}\penalty 10000\ \mathcal{L}_{reg}+\alpha_{3}\penalty 10000\ \mathcal{L}_{cs}$ (8)
where, $\alpha_{1},\alpha_{2},\alpha_{3}$ are tunable weights summing to 1, controlling the influence of each component.
The classification phase refines prototypes into interpretable classes, enhancing clustering and enabling generalizable, multi-granular predictions.
The computational complexity of ProtoSiTex is analyzed in Appendix A of the supplementary file. Appendix B provides a detailed discussion on the stability of prototype updates and the influence of loss-weight parameters on model behavior.
IV Experiments
This section describes the datasets employed and reports the corresponding experimental results.
IV-A Datasets
To evaluate the effectiveness of ProtoSiTex and compare it against state-of-the-art (SOTA) methods, we conducted comprehensive experiments on benchmark datasets IMDb [IMDb], TweetEVAL [TweetEVAL], and our newly proposed hotel reviews dataset (HR). These datasets encompass a diverse range of classification tasks: binary, multi-class, and multi-label, respectively.
(i) IMDb [IMDb]: Consisting of 50K movie reviews for binary sentiment classification (positive/ negative).
(ii) TweetEVAL [TweetEVAL]: A multiclass dataset comprising 5052 tweets, with each tweet annotated with one of four emotion labels: anger, joy, optimism, or sadness.
07507501,5001{,}5002,2502{,}2503,0003{,} [^0] Room & AmenitiesPricing & ValuePersonalizationLocation & SurroundingHousekeepingGuest ExperienceFoodConcierge2,1872{,}1879009009769761,9551{,}9551,8501{,}8502,4172{,}4172,2852{,}2852,2552{,}255Document-level review count
Figure 2: Distribution of eight aspects in HR
(iii) Proposed Dataset on Hotel Reviews (HR): To support fine-grained, interpretable, multi-aspect classification in the hospitality domain, we introduce a novel dataset comprising hotel reviews, curated from tripadvisor.com. The reviews were collected via web scraping, ensuring a diverse coverage across locations, and customer sentiments. We identified $8$ key aspects based on extensive consultations with experts in travel management and linguistics (Fig. 2). These aspects were chosen to reflect both operational and experiential dimensions of hotel service quality. Each review may correspond to multiple aspect labels, reflecting the inherently multi-label nature of user opinions. To ensure high-quality annotations, we employed $10$ trained human annotators for each review. For every review, the annotators assigned one or more aspect labels as applicable, and highlighted the exact subsentence or phrase responsible for each aspect label, thereby providing ground-truth supervision for interpretability evaluation. Prior to annotation, all annotators underwent training. We implemented a quality control mechanism using in-context learning via GPT-4o [ARTICLE_AAAI_2025]. Annotator performance was iteratively assessed by comparing human-labeled segments, enabling correction and refinement through active feedback loops. To ensure annotation consistency and reliability, we quantified inter-annotator agreement using Krippendorff’s alpha ($\upalpha$) [krippendorff1970]. The obtained $\upalpha=0.85$ indicates strong inter-annotator agreement, validating the annotation procedure. The resulting dataset consists of 3002 document-level and 26519 sentence-level reviews, annotated at the subsentence level with multi-label aspect tags and text spans responsible for each label, making it one of the first of its kind to support interpretability-grounded multi-aspect classification at such fine granularity. Brief statistical details of HR are presented in Table I.
Notably, subsentence level annotations are absent in the IMDb and TweetEVAL datasets.
IV-B Experimental Setup
All experiments were conducted on a Dell Precision 5860 Tower equipped with PyTorch 2.5.1, CUDA 12.2, an NVIDIA RTX A6000 GPU with 48 GB VRAM, and 256 GB RAM. The HR dataset was split into training, validation, and test sets in 7:2:17{:}2{:}1. For IMDb [IMDb] and TweetEVAL [TweetEVAL], we used the given training/ validation/ test sets. Each experiment was run five times, and the mean test performance is reported here.
Table II presents the hyperparameters engaged in ProtoSiTex. To obtain high-dimensional, contextually rich text embeddings, we incorporate a language model, called RoBERTa [liu2019roberta]. Specifically, it is a robustly optimized BERT variant, which we employed in its large-weight version pretrained on large English-language internet corpora using the masked language modeling objective; RoBERTa-Large remains tunable within our framework. We maintain a consistent dropout rate of 0.1 across all trainable weights. To avoid overfitting, early stopping with a patience of three epochs based on validation performance is applied. Additionally, we set a maximum training limit of 100 epochs. Each epoch includes 10 ($i_{1}$) iterations of the clustering phase and 100 ($i_{2}$) iterations of the classification phase, ensuring a balanced optimization of both phases.
TABLE I: Statistics of proposed dataset: Hotel Reviews (HR)
Attributes Values No. of Hotels & Locations 275 & 70 No. of Reviews & Sentences 3002 & 27074 No. of Aspects 8 Sentences per review (min & max) 1 & 92 Sentences per review (mean $\pm$ stdev) 9.01 $\pm$ 7.69 Words per review (mean) 120.82
TABLE II: Brief details of hyperparameters
Attribute Value Optimizer AdamW Learning rate (Clustering-phase) $2\times 10^{-5}$ Learning rate (Classification-phase) $2\times 10^{-5}$ Batch size 64 LM encoder: Feature dimension 1024 No. of dense layer 2 Dimension of dense layer 256 & 64 Clustering loss parameters ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ ) [0-1] Classification loss parameters ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$) [0-1] No. of attention heads ($h$) [1-32] No. of learnable prototypes ($q$) [4-64]
TABLE III: Comparison with baseline and SOTA approaches
[…]
IV-C Evaluation Metrics
We evaluate model performance using task-specific metrics. For binary and multi-class classification, we report Accuracy ($\mathcal{A}$) and F1-score ($\mathcal{F}$); aspect-wise evaluation additionally includes Precision ($\mathcal{P}r$) and Recall ($\mathcal{R}$); for multi-label classification, we use macro-averaged F1-score ($\mathcal{F}_{m}$) and Balanced Accuracy ($\mathcal{BA}_{m}$); all are in percentage. We also compute Hamming loss ($\mathcal{HL}$) to analyze label-wise prediction performance, and report complementary Hamming loss ${\mathcal{HL}}^{c}$ here; ${\mathcal{HL}}^{c}=(1-\mathcal{HL})\times 100\%$.
IV-D Results & Comparative Study
To compare with ProtoSiTex, we chose 6 SOTA prototypical learning models: ProSeNet [10.1145/3292500.3330908-ProsNet], ProtoTEX [das-etal-2022-prototex], ProtoryNet [10.5555/3648699.3648963-ProtoryNet], Proto-lm [xie-etal-2023-proto], GAProtoNet [wen-etal-2025-gaprotonet], and ProtoLens [protolens]. In addition to this, we selected 9 black-box LMs: ALBERT [ALBERT], DeBERTaV3 [DeBERTaV3] BERT [devlin-etal-2019-bert], DistilBERT [sanh2019distilbert], ELECTRA [ELECTRA], XLNET [yang2019xlnet], ModernBERT [ModernBERT], RoBERTa and RoBERTa-Large [liu2019roberta], and 2 text classification methods: TextGCN [textgcn] and BertGCN [bertgcn], without incorporating interpretability.
In Table III, we present the comparative study. This study reveals that existing prototype-based SOTA approaches often sacrifice performance for interpretability, leading to significantly lower results compared to baselines, particularly on challenging multi-label datasets such as HR. In contrast, ProtoSiTex successfully overcame this limitation by narrowing the performance gap with black-box LMs, while preserving inherent interpretability across varying dataset complexities. Specifically, ProtoSiTex achieved improvements of 0.04% and 2.85% in $\mathcal{A}$} on the IMDb and TweetEVAL datasets, respectively, and delivered a substantial 9.98 % improvement in $\mathcal{F}_{m}$} on HR, compared to the best prior prototype-based methods. These results across three diverse tasks highlight ProtoSiTex’s generalizability and applicability.
${\mathcal{F}}_{m}$${\mathcal{BA}}_{m}$${\mathcal{HL}}^{c}$75758080858590909595K-meansKaimingXavierRandom
${\mathcal{F}}_{m}$${\mathcal{BA}}_{m}$${\mathcal{HL}}^{c}$6060707080809090100100w/o ATProtoSiTex
${\mathcal{F}}_{m}$${\mathcal{BA}}_{m}$${\mathcal{HL}}^{c}$2020404060608080100100Euclidean Dist.Cosine Sim.ProtoSiTex
[…]
Figure 3: Performance analysis of ProtoSiTex on HR: (a) with various prototype initializers; (b) without alternate training (w/o AT); (c) across different similarity measures; with a varying number of (d) prototypes, and (e) heads
[…]
Figure 4: Impact of hyperparameters $\alpha_{i}$ and $\lambda_{i}$ on HR
IV-E Impact of Model Components
We here report the impact of modules of ProtoSiTex on HR dataset only due to space constraints, which is representative enough to illustrate the observed trends.
IV-E1 Impact of Different Prototype Initializers
We evaluated four prototype initialization strategies: random, K-means centroids, Kaiming, and Xavier. As shown in Fig. 3(a), random initialization consistently outperformed the others across $\mathcal{F}_{m}$, $\mathcal{BA}_{m}$, and ${\mathcal{HL}}^{c}$. This performance gain may be attributed to the increased diversity introduced early in the training, which helps avoid premature convergence and overfitting.
IV-E2 Impact of Alternate Training
ProtoSiTex adopts a dual-phase training strategy, alternating between unsupervised prototype learning and supervised classification. To evaluate its efficacy, we compared model variants trained with and without alternation. As shown in Fig. 3(b), alternate training significantly improved performance by enabling the model to discover semantically meaningful prototype clusters, and align them with downstream classification objectives. In contrast, without alternation, the model struggled to learn task-specific prototypes, lowering performance.
IV-E3 Impact of Similarity Function
We evaluated three similarity metrics: MHA, cosine similarity, and Euclidean distance, for associating subsentence embeddings with prototypes. As presented in Fig. 3(c), MHA of ProtoSiTex achieved superior $\mathcal{F}_{m}$, $\mathcal{BA}_{m}$, and ${\mathcal{HL}}^{c}$ scores by capturing richer embedding relationships, leading to more discriminative and interpretable prototype assignments.
IV-F Impact of Hyperparameters
ProtoSiTex includes several key hyperparameters: the number of prototypes ($q$), number of heads ($h$), loss weights for clustering ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$), and classification ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$). To systematically investigate the influence of these hyperparameters on model performance, we conducted a series of experiments on the HR. As shown in Fig. 3(d), increasing $q$ initially boosts performance by capturing diverse semantic concepts, but excessive values lead to redundancy and over-smoothing. A similar trend is observed for $h$ in Fig. 3(e). Furthermore, Fig. 4 shows that assigning $\alpha_{1}=0.01$, $\alpha_{2}=0.98$, $\alpha_{3}=0.01$, and $\lambda_{1}=0.6$, $\lambda_{2}=0.2$, $\lambda_{3}=0.2$ yields optimal results. This multi-level supervision prevented the model from overfitting to a single level of granularity, supporting both local and global interpretability and accuracy.
[…]
Figure 5: Aspect-wise performance of ProtoSiTex on HR
(a) Refer to caption
(b) Refer to caption (c) Refer to caption
Figure 6: Correctly predicted examples of ProtoSiTex on (a) HR, (b) TweetEVAL [TweetEVAL], and (c) IMDb [IMDb]. Best viewed in color.
(a) Refer to caption
(b) Refer to caption (c) Refer to caption
Figure 7: Misprediction examples of ProtoSiTex on (a) HR, (b) TweetEVAL [TweetEVAL], and (c) IMDb [IMDb]. Best viewed in color.
IV-G Aspect-wise Performance Analysis
Fig. 5 presents a detailed aspect-wise evaluation of ProtoSiTex on the HR dataset, offering insights into the model’s performance across individual aspect categories. Notably, the aspects Guest Experience, Food, and Concierge not only appear the most frequently in HR but also exhibit strong predictive performance. In contrast, Pricing & Value and Personalization are the least represented. Despite the challenges posed by class imbalance, ProtoSiTex shows a strong ability to identify the Pricing & Value aspect, achieving $\mathcal{F}$ and $\mathcal{A}$ scores comparable to those of the more frequent categories. However, Personalization remains the most challenging aspect, particularly in terms of $\mathcal{R}$, as the model often fails to capture all relevant instances. This difficulty is likely due to both its low frequency in the dataset and its inherent subjective nature. Unlike more explicit aspects such as Food or Location & Surrounding, Personalization is often expressed through subtle, indirect, and context-dependent language, making it more difficult to detect.
IV-H Qualitative Analysis of ProtoSiTex
To illustrate ProtoSiTex’s interpretability and prediction quality, we present correctly predicted examples from the HR, TweetEVAL, and IMDb datasets in Fig. 6. In each case here, the model accurately predicts aspect or sentiment labels while grounding them in semantically coherent subsentence segments via learned prototypes. In the HR example, ProtoSiTex correctly identifies multi-label aspects (Room & Amenities, Food, Pricing & Value, etc.) from a single review, while providing text-grounded justifications. This highlights its ability to disentangle overlapping review aspects and attribute them to distinct prototype-based explanations. In the TweetEVAL sample, ProtoSiTex captures the emotion joy and connects it to meaningful prototypes reflecting positive sentiment, demonstrating its effectiveness in emotional reasoning with minimal text context. Similarly, in the IMDb example, ProtoSiTex predicts the positive label and supports it with a rich prototype cluster centered on artistic and narrative appreciation, confirming alignment with human sentiment. Overall, these examples demonstrate that ProtoSiTex offers not only accurate predictions but also transparent, faithful interpretations at the subsentence level, making it suitable for real-world applications requiring interpretability.
IV-I Misprediction Analysis
To better understand the behavior and limitations of ProtoSiTex, we conducted a detailed analysis of mispredictions. This analysis highlights recurring patterns that lead to incorrect predictions. The most notable cases of mispredictions are discussed below:
IV-I1 Boundary Ambiguity in Prototype
In multi-label classification, ProtoSiTex occasionally exhibits boundary ambiguity among prototypes, where semantically related class representations capture overlapping evidence. This ambiguity leads to confusion in distinguishing co-occurring categories, resulting in the assignment of only a single dominant label instead of multiple relevant ones. For example, in Fig. 7(a), the review “Spacious room with balcony with amazing view of the valley and the falls”. The actual labels are Room & Amenities and Location & Surrounding. However, ProtoSiTex predicts only Room & Amenities. Here, Room & Amenities dominates and captures shared contextual characteristics such as “spacious room”, “balcony”, overshadowing the equally valid cues of “view of the valley and the falls” that should activate Location & Surrounding.
IV-I2 Lexical Cue Dominance
Sometimes the mispredictions in ProtoSiTex were caused by lexical bias, the tendency to rely excessively on high-intensity words when making predictions. Words such as “angry”, “worst”, or “best” often dominate the model’s decision process, even when their contextual meaning is softened/ inverted by surrounding text. For instance, in Fig. 7(b) the review “@user i love ur angry comments”, the actual emotion is affectionate teasing or joy, yet the ProtoSiTex incorrectly predicts anger due to the presence of the word “angry”.
IV-I3 Absence of Fine-Grained Annotation
This observation highlights an important takeaway: even when only coarse document-level supervision is available, prototype-based explanations remain highly useful for error analysis and model trust. By showing how individual subsentences align with prototypes tied to different target classes, ProtoSiTex offers interpretable evidence that pinpoints why a misclassification occurred, rather than treating it as a black-box error. From a technical perspective, the misclassification arises due to aggregation bias, while subsentence–prototype mappings are accurate, the hierarchical aggregation process (from prototype-level activations to the final document-level decision) can overweight strong local cues and misjudge the overall label. In practice, this means even when the document-level prediction is wrong, the prototype-level reasoning still faithfully explains the decision path, helping to uncover issues such as prototype imbalance, lack of discourse-level cues, or dataset limitations tied to coarse supervision. Thus, beyond accuracy metrics, our approach demonstrates that interpretable prototype assignments provide actionable insights for debugging and model refinement. As illustrated in Fig. 7(c), the identified prototypes, explanations, and associated class mappings are correct, even though the final document-level prediction was misclassified.
Additional experimental analyses are provided in the supplementary file. Appendix C presents the convergence behavior of the individual and joint loss functions during training. Appendix D offers a comprehensive evaluation of prototype interpretability using multiple metrics, including coverage, contrastivity, centered kernel alignment, and fidelity. Appendix E reports a detailed nonparametric statistical significance analysis based on the Friedman and Nemenyi tests. Finally, Appendix F provides extensive qualitative results, illustrating representative prototypes generated across the three benchmark datasets.
V Conclusion
We proposed ProtoSiTex, a semi-interpretable framework for fine-grained, multi-label text classification. By combining adaptive prototype learning, dual-phase alternate training, and hierarchical supervision, ProtoSiTex delivers accurate and transparent predictions across multiple levels of granularity. To support evaluation, we introduced the hotel reviews dataset (HR) with subsentence-level multi-aspect annotations. Extensive experiments on HR, along with the publicly available IMDb and TweetEVAL datasets, show that ProtoSiTex matches or exceeds baseline and state-of-the-art models, while providing faithful, human-aligned interpretations. Future directions include extending ProtoSiTex to cross-domain and multilingual settings, incorporating multimodal cues, and integrating large language models for interactive, user-centric explanations.
Acknowledgment
The authors heartily thank all the researchers/ interns/ annotators who helped in hotel reviews dataset preparation.